a truck P (900kg)starts from rest runs down a frictionless slope from point a to b which is 101.4 meters, from a height at point a of 33.8 meters to b 0 meters,truck P collided with truck Q (337,5kg)at point b and continues to move with a speed of 19,5 m.s1.,what is tuck Q velocity after collision


1
Expert's answer
2021-03-02T18:16:12-0500

We have,


Mass of truck P, "m_1" = 900 kg


Mass of truck Q, "m_2" = 337.5kg


As the initial velocity of truck P = 0 m/s


Using the third equation of motion,


"v^2 - u^2 = 2\times gcos\theta\times s"


"v^2 = 2\times 10 \times 0.33\times 101.4"


"v = 25.86" m/s


Now using the conservation of momentum,


"m_1u_1 + m_2u_2 = m_1v_1 + m_2v_2" ( Here "u_1 = v" )


"900\times 25.86 + 0 = 900\times 19.5 + 337.5\times v_2"


"v_2 = \dfrac {5724}{337.5}"


"v_2 = 16.96 m\/s"

Hence , the velocity of truck Q after collision is "16.96" m/s